# Kengo Nakamura, Masaaki Nishino, and Shuhei Denzumi NTT Communication Science Laboratories, Kyoto, Japan

### v1

@@ 3.2.3 Restrict, Permit, Nonsuperset, and Nonsubset
Next, we consider the restrict, permit, nonsuperset, and nonsubset operations. These operations include inclusion relations of subsets in their definitions, which makes it difficult to generate a hidden weighted bit function as a result of the operation. This is due to the fact that $\mathcal{H}_{m}$ includes the universal set $Y$ as well as a singleton $\{y_{1}\}$. For example, if $\mathcal{F}$ is the family of subsets of $Y$ and the universal set $Y$ is included in the result of $\mathcal{F}\oslash\mathcal{G}$, all of the subsets in $\mathcal{F}$ must be included in $\mathcal{F}\oslash\mathcal{G}$ thanks to the definition of the restrict operation.
Instead, we use the permutation function. Because all of the subsets in $\mathcal{P}_{m}$ have cardinality $m$, the above issue can be alleviated. More specifically, we prepared the complement of the families:
$\mathcal{C}_{m}$$\coloneqq\{S\subseteq\{y_{1},\ldots,y_{m^{2}}\}\mid|S|=m\},$
$\mathcal{T}_{m,k}$ $\coloneqq\mathcal{C}_{m}\setminus\mathcal{Q}_{m,k}(=\mathcal{C}_{m}\cap(2^{Y}\setminus\mathcal{Q}_{m,k})).$
@@
$\bigcup_{k=1}^{2m}\mathcal{T}_{m,k}$ $=\mathcal{C}_{m}\cap\left(\bigcup_{k=1}^{2m}(2^{Y}\setminus\mathcal{Q}_{m,k})\right)$
$=\mathcal{C}_{m}\cap\left(2^{Y}\setminus\left(\bigcap_{k=1}^{2m}\mathcal{Q}_{m,k}\right)\right)$
$=\mathcal{C}_{m}\cap(2^{Y}\setminus\mathcal{P}_{m})=\mathcal{C}_{m}\setminus\mathcal{P}_{m}.$
We use these families $\mathcal{T}_{m,k}$ to prove the theorems.
Theorem 9.
@@
Figure 4: Example of blow-up for permit operation.
Proof.
@@
As proved in Appendix, $Z(\mathcal{C}_{m})=O(m^{3})$ and $Z(\mathcal{T}_{m,k})=O(m^{3})$. Thus, $Z(\mathcal{F}_{m})=O(m^{3})$ and $Z(\mathcal{G}_{m})=O(m^{4})$. Any set in $\mathcal{F}_{m}=\mathcal{C}_{m}$ consists of $m$ elements chosen from $y_{1},\ldots,y_{m^{2}}$, and any set in $\mathcal{G}_{m}$ consists of $m$ elements from $y_{1},\ldots,y_{m^{2}}$ plus one element from $x_{1},\ldots,x_{2m}$. Thus, set $S\in\mathcal{F}_{m}$ is a subset of some sets in $\mathcal{G}_{m}$ if and only if $\{x_{k}\}\cup S\in\mathcal{G}_{m}$ for some $k$. In other words, $S\in\mathcal{F}\oslash\mathcal{G}$ if and only if $S$ is included in $\mathcal{T}_{m,k}$ for some $k$. Since $\mathcal{C}_{m}\supset\mathcal{T}_{m,k}$ for any $k$ by definition, this means $\mathcal{F}_{m}\oslash\mathcal{G}_{m}=\bigcup_{k=1}^{2m}\mathcal{T}_{m,k}=\mathcal{C}_{m}\setminus\mathcal{P}_{m}$. Since $Z(\mathcal{C}_{m})=O(m^{3})$ and $Z(\mathcal{P}_{m})=\Omega(2^{m}/m)$, we have $Z(\mathcal{F}_{m}\oslash\mathcal{G}_{m})=\Omega(2^{m}/m^{4})$ by Lemma 7. The ZDDs involved are depicted in Figure 4.
@@
The restrict and nonsuperset operations can be handled by nearly the same families. We define the same $\mathcal{G}_{m}$ and let $\mathcal{F}_{m}\coloneqq\{X\}\sqcup\mathcal{C}_{m}$. Similar to the proof of the permit operation, set $X\cup S\in\mathcal{F}_{m}$ ($S\subseteq Y$) is a superset of some sets in $\mathcal{G}_{m}$ if and only if $\{x_{k}\}\cup S\in\mathcal{G}_{m}$ for some $k$. This means $\mathcal{F}_{m}\bigtriangleup\mathcal{G}_{m}=\{X\}\sqcup(\bigcup_{k=1}^{2m}\mathcal{T}_{m,k})=\{X\}\sqcup(\mathcal{C}_{m}\setminus\mathcal{P}_{m})$, whose ZDD size is $\Omega(2^{m}/m^{4})$. For the nonsuperset operation, we have $\mathcal{F}_{m}\operatorname{\searrow}\mathcal{G}_{m}=\mathcal{F}_{m}\setminus(\mathcal{F}_{m}\bigtriangleup\mathcal{G}_{m})=\{X\}\sqcup\mathcal{P}_{m}$, yielding $Z(\mathcal{F}_{m}\operatorname{\searrow}\mathcal{G}_{m})=\Omega(2^{m}/m)$. ∎
@@ 3.2.4 Maximal and Minimal
For the maximal and minimal operations, we use the close relationship with the nonsuperset and nonsubset operations. We prepare a family having $\mathcal{F}_{m}$ and $\mathcal{G}_{m}$ appearing in the proof of Theorem 9 as a subfamily.
Theorem 10.
Let ${}^{\diamond}$ be a unary operator chosen from maximal (${}^{\uparrow}$) and minimal (${}^{\downarrow}$). Then, there exists a sequence of family $\mathcal{F}_{m}$ such that (i) $\mathcal{F}_{m}$ is a family of subsets of a set of $O(m^{2})$ elements, (ii) $Z(\mathcal{F}_{m})=O(m^{4})$, and (iii) $Z(\mathcal{F}_{m}^{\diamond})=\Omega(2^{m}/\mathrm{poly}(m))$.
Figure 5: Example of blow-up for maximal operation.
Proof.
@@
$\mathcal{F}_{m}$ $\coloneqq\mathcal{C}_{m}\cup\left[\{\{w\}\}\sqcup\mathcal{G}_{m}\right],$
$\text{where}\quad\mathcal{G}_{m}$ $\coloneqq\bigcup_{k=1}^{2m}(\{\{x_{k}\}\}\sqcup\mathcal{T}_{m,k}).$
Here, we observe that this $\mathcal{G}_{m}$ is the same as that appearing in the proof of Theorem 9. The ZDD size is bounded as $Z(\mathcal{F}_{m})=O(Z(\mathcal{C}_{m})+Z(\mathcal{G}_{m}))=O(m^{4})$. Every set in $\mathcal{C}_{m}$ has $m$ elements and every set in $\{\{w\}\}\sqcup\mathcal{G}_{m}$ has $m+2$ elements. Thus, every set in the latter family is maximal, while a set in the former family is maximal if and only if it is not a subset of any set included in the latter family. Therefore, we have
$\mathcal{F}_{m}^{\uparrow}$ $=\left[\mathcal{C}_{m}\operatorname{\nearrow}(\{\{w\}\}\sqcup\mathcal{G}_{m})\right]\cup\left[\{\{w\}\}\sqcup\mathcal{G}_{m}\right]$
$=\left[\mathcal{C}_{m}\operatorname{\nearrow}\mathcal{G}_{m}\right]\cup\left[\{\{w\}\}\sqcup\mathcal{G}_{m}\right]$
$=\mathcal{P}_{m}\cup\left[\{\{w\}\}\sqcup\mathcal{G}_{m}\right],$
where the second equality holds because all of the sets in $\mathcal{C}_{m}$ do not include $w$ and the last equality follows from the proof of Theorem 9. The resultant ZDD is like the right one in Figure 5, which implies $Z(\mathcal{F}_{m}^{\uparrow})\geq Z(\mathcal{P}_{m})=O(2^{m}/m)$.
The minimal operation can be treated in a similar way. We define $\mathcal{F}_{m}$ as
@@
where $\mathcal{G}_{m}$ is the same family as that above. We again have $Z(\mathcal{F}_{m})=O(m^{4})$. Every set in $\mathcal{G}_{m}$ has $m+1$ elements and every set in $\{\{w\}\}\sqcup\{\{x_{1},\ldots,x_{2m}\}\}\sqcup\mathcal{C}_{m}$ has $13m+$ 1 elements. Thus, every set in the former family is minimal, while a set in the latter family is minimal if and only if it is not a superset of any set included in the former family. Now we have
$\mathcal{F}_{m}^{\downarrow}$ $=\mathcal{G}_{m}\cup\left[(\{\{w\}\}\sqcup\{\{x_{1},\ldots,x_{2m}\}\}\sqcup\mathcal{C}_{m})\operatorname{\searrow}\mathcal{G}_{m}\right]$
$=\mathcal{G}_{m}\cup\left[(\{\{w\}\}\sqcup\{\{x_{1},\ldots,x_{2m}\}\}\sqcup\mathcal{C}_{m})\operatorname{\searrow}\mathcal{G}_{m}\right]$
$=\mathcal{G}_{m}\cup\left[\{\{w\}\}\sqcup\{\{x_{1},\ldots,x_{2m}\}\}\sqcup\mathcal{P}_{m}\right],$
@@
where the second equality holds because none of the sets in $\mathcal{G}_{m}$ includes $w$ and the last equality follows from the proof of Theorem 9. This again implies $Z(\mathcal{F}_{m}^{\downarrow})\geq Z(\mathcal{P}_{m})=O(2^{m}/m)$. ∎
3.2.5 Minimal Hitting Set
Finally, we prove the minimal hitting set operation. We can prove that a simple family’s minimal hitting set becomes the permutation function.
Theorem 11.
@@ 3.2.5 Minimal Hitting Set
There exists a sequence of family $\mathcal{F}_{m}$ such that (i) $\mathcal{F}_{m}$ is a family of subsets of a set of $O(m^{2})$ elements, (ii) $Z(\mathcal{F}_{m})=O(m^{2})$, and (iii) $Z(\mathcal{F}_{m}^{\sharp})=\Omega(2^{m}/m)$.
Proof.
We consider a family of subsets of $\{y_{1},\ldots,y_{m^{2}}\}$. For $k=1,\ldots,m$, we set $S_{k}\coloneqq\{y_{m(k\!-\!1)\!+\!1},y_{m(k\!-\!1)\!+\!2},\ldots,y_{m(k\!-\!1)\!+\!m}\}$, and $S_{m+k}\coloneqq\{y_{k},y_{m\!+\!k},\ldots,y_{m(m\!-\!1)\!+\!k}\}$. Using them, we define $\mathcal{F}_{m}\coloneqq\{S_{1},\ldots,S_{2m}\}$. Then, we can easily derive $Z(\mathcal{F}_{m})=O(m^{2})$.
@@
For $S\subseteq\{y_{1},\ldots,y_{m^{2}}\}$, we associate a binary $m\times m$ matrix, where the $(i,j)$-element is $1$ if and only if $y_{m(i-1)+j}\in S$. Then, $S\cap S_{k}\neq\emptyset$ means that the $k$-th row of the matrix has at least one $1$ and $S\cap S_{m+k}\neq\emptyset$ means that the $k$-th column of the matrix has at least one $1$. Thus, $S\in\mathcal{F}_{m}^{\sharp}$ if and only if the corresponding matrix has at least one $1$ for any column or row and no proper subset of $S$ satisfies this property. The minimal matrix having this property is the permutation matrix, and thus $\mathcal{F}_{m}^{\sharp}=\mathcal{P}_{m}$, that is, the permutation function. This implies $Z(\mathcal{F}_{m}^{\sharp})=\Omega(2^{m}/m)$. ∎

### v2

@@ 3.2.3 Restrict, Permit, Nonsuperset, and Nonsubset
These operations include inclusion relations of subsets in their definitions, which makes it difficult to generate a hidden weighted bit function as a result of the operation. This is due to the fact that $\mathcal{H}_{m}$ includes the universal set $Y$ as well as a singleton $\{y_{1}\}$. For example, if $\mathcal{F}$ is the family of subsets of $Y$ and the universal set $Y$ is included in the result of $\mathcal{F}\oslash\mathcal{G}$, all of the subsets in $\mathcal{F}$ must be included in $\mathcal{F}\oslash\mathcal{G}$ due to the definition of the permit operation.
Instead, we use the permutation function. Because every set in $\mathcal{P}_{m}$ has cardinality $m$, the above issue can be alleviated. More specifically, we prepared the complement of the families:
$\mathcal{C}_{m}\coloneqq\{S\subseteq\{y_{1},\ldots,y_{m^{2}}\}\mid|S|=m\},\quad\mathcal{T}_{m,k}\coloneqq\mathcal{C}_{m}\setminus\mathcal{Q}_{m,k}(=\mathcal{C}_{m}\cap(2^{Y}\setminus\mathcal{Q}_{m,k})).$
@@
$\bigcup_{k=1}^{2m}\mathcal{T}_{m,k}$ $=\mathcal{C}_{m}\cap\left(\bigcup_{k=1}^{2m}(2^{Y}\setminus\mathcal{Q}_{m,k})\right)=\mathcal{C}_{m}\cap\left(2^{Y}\setminus\left(\bigcap_{k=1}^{2m}\mathcal{Q}_{m,k}\right)\right)=\mathcal{C}_{m}\setminus\mathcal{P}_{m}.$
We use these families $\mathcal{T}_{m,k}$ to prove the following.
Theorem 3.11.
@@
Figure 3: Example of blow-up for permit (left) and maximal (right) operations.
Proof 3.12.
@@
As proved in Section 3.4, $Z(\mathcal{C}_{m})=O(m^{3})$ and $Z(\mathcal{T}_{m,k})=O(m^{3})$. Thus, $Z(\mathcal{F}_{m})=O(m^{3})$ and $Z(\mathcal{G}_{m})=O(m^{4})$. Any set in $\mathcal{F}_{m}=\mathcal{C}_{m}$ consists of $m$ elements chosen from $y_{1},\ldots,y_{m^{2}}$, and any set in $\mathcal{G}_{m}$ consists of $m$ elements from $y_{1},\ldots,y_{m^{2}}$ plus one element from $x_{1},\ldots,x_{2m}$. Thus, set $S\in\mathcal{F}_{m}$ is a subset of some set in $\mathcal{G}_{m}$ if and only if $\{x_{k}\}\cup S\in\mathcal{G}_{m}$ for some $k$. In other words, $S\in\mathcal{F}_{m}\oslash\mathcal{G}_{m}$ if and only if $S$ is included in $\mathcal{T}_{m,k}$ for some $k$. Since $\mathcal{C}_{m}\supset\mathcal{T}_{m,k}$ for any $k$ by definition, this means $\mathcal{F}_{m}\oslash\mathcal{G}_{m}=\bigcup_{k=1}^{2m}\mathcal{T}_{m,k}=\mathcal{C}_{m}\setminus\mathcal{P}_{m}$. Since $Z(\mathcal{C}_{m})=O(m^{3})$ and $Z(\mathcal{P}_{m})=\Omega(2^{m}/m)$, we have $Z(\mathcal{F}_{m}\oslash\mathcal{G}_{m})=\Omega(2^{m}/m^{4})$ by Lemma 3.7. The ZDDs involved are depicted in Figure 3.
@@
The restrict and nonsuperset operations can be handled by nearly the same families. We define the same $\mathcal{G}_{m}$ and let $\mathcal{F}_{m}\coloneqq\{X\}\sqcup\mathcal{C}_{m}$. Similar to the proof of the permit operation, set $X\cup S\in\mathcal{F}_{m}$ ($S\subseteq Y$) is a superset of some sets in $\mathcal{G}_{m}$ if and only if $\{x_{k}\}\cup S\in\mathcal{G}_{m}$ for some $k$. This means $\mathcal{F}_{m}\bigtriangleup\mathcal{G}_{m}=\{X\}\sqcup(\bigcup_{k=1}^{2m}\mathcal{T}_{m,k})=\{X\}\sqcup(\mathcal{C}_{m}\setminus\mathcal{P}_{m})$, whose ZDD size is $\Omega(2^{m}/m^{4})$. For the nonsuperset operation, we have $\mathcal{F}_{m}\operatorname{\searrow}\mathcal{G}_{m}=\mathcal{F}_{m}\setminus(\mathcal{F}_{m}\bigtriangleup\mathcal{G}_{m})=\{X\}\sqcup\mathcal{P}_{m}$, yielding $Z(\mathcal{F}_{m}\operatorname{\searrow}\mathcal{G}_{m})=\Omega(2^{m}/m)$.
@@ 3.2.4 Maximal and Minimal
For these operations, we use the close relationship with the nonsuperset and nonsubset operations. We prepare a family having $\mathcal{F}_{m}$ and $\mathcal{G}_{m}$ appearing in the proof of Theorem 3.11 as a subfamily.
Theorem 3.13.
Let be a unary operator chosen from maximal () and minimal (). Then, there exists a sequence of families $\mathcal{F}_{m}$ such that (i) $\mathcal{F}_{m}$ is a family of subsets of a set of $O(m^{2})$ elements, (ii) $Z(\mathcal{F}_{m})=O(m^{4})$, and (iii) $Z(\mathcal{F}_{m}^{\diamond})=\Omega(2^{m}/\mathrm{poly}(m))$.
Proof 3.14.
@@
$\mathcal{F}_{m}\coloneqq\mathcal{C}_{m}\cup\left[\{\{w\}\}\sqcup\mathcal{G}_{m}\right],\quad\text{where}\ \mathcal{G}_{m}\coloneqq\bigcup_{k=1}^{2m}(\{\{x_{k}\}\}\sqcup\mathcal{T}_{m,k}).$
Here, we observe that this $\mathcal{G}_{m}$ is the same as that appearing in the proof of Theorem 3.11. The ZDD size is bounded as $Z(\mathcal{F}_{m})=O(Z(\mathcal{C}_{m})+Z(\mathcal{G}_{m}))=O(m^{4})$. Every set in $\mathcal{C}_{m}$ has $m$ elements and every set in $\{\{w\}\}\sqcup\mathcal{G}_{m}$ has $m+2$ elements. Thus, every set in the latter family is maximal, while a set in the former family is maximal if and only if it is not a subset of any set included in the latter family. Therefore, we have
$\mathcal{F}_{m}^{\uparrow}$ $=\left[\mathcal{C}_{m}\operatorname{\nearrow}(\{\!\{\!w\!\}\!\}\sqcup\mathcal{G}_{m})\right]\!\cup\!\left[\{\!\{\!w\!\}\!\}\sqcup\mathcal{G}_{m}\right]=\left[\mathcal{C}_{m}\operatorname{\nearrow}\mathcal{G}_{m}\right]\!\cup\!\left[\{\!\{\!w\!\}\!\}\sqcup\mathcal{G}_{m}\right]=\mathcal{P}_{m}\cup\left[\{\!\{\!w\!\}\!\}\sqcup\mathcal{G}_{m}\right],$
where the second equality holds because all of the sets in $\mathcal{C}_{m}$ do not include $w$ and the last equality follows from the proof of Theorem 3.11. The resultant ZDD is like the right one in Figure 3, which implies $Z(\mathcal{F}_{m}^{\uparrow})\geq Z(\mathcal{P}_{m})=\Omega(2^{m}/m)$.
The minimal can be treated in a similar way. We define
@@
where $\mathcal{G}_{m}$ is the same family as that above. We again have $Z(\mathcal{F}_{m})=O(m^{4})$. Every set in $\mathcal{G}_{m}$ has $m+1$ elements and every set in $\{\{w\}\}\sqcup\{X\}\sqcup\mathcal{C}_{m}$ has $13m+$ 1 elements. Thus, every set in the former family is minimal, while a set in the latter family is minimal if and only if it is not a superset of any set included in the former family. Now we have
$\mathcal{F}_{m}^{\downarrow}$ $=\mathcal{G}_{m}\cup\left[(\{\{w\}\}\sqcup\{X\}\sqcup\mathcal{C}_{m})\operatorname{\searrow}\mathcal{G}_{m}\right]$
$=\mathcal{G}_{m}\cup\left[\{\{w\}\}\sqcup((\{X\}\sqcup\mathcal{C}_{m})\operatorname{\searrow}\mathcal{G}_{m})\right]=\mathcal{G}_{m}\cup\left[\{\{w\}\}\sqcup\{X\}\sqcup\mathcal{P}_{m}\right],$
@@
where the second equality holds because none of the sets in $\mathcal{G}_{m}$ includes $w$ and the last equality follows from the proof of Theorem 3.11. This again implies $Z(\mathcal{F}_{m}^{\downarrow})\geq Z(\mathcal{P}_{m})=\Omega(2^{m}/m)$.
3.2.5 Minimal Hitting Set and Closure
For these operations, we can constitute much simpler examples.
Theorem 3.15.
@@ 3.2.5 Minimal Hitting Set and Closure
Let be a unary operator chosen from minimal hitting set () and closure (). Then, there exists a sequence of families $\mathcal{F}_{m}$ such that (i) $\mathcal{F}_{m}$ is a family of subsets of a set of $O(m^{2})$ elements, (ii) $Z(\mathcal{F}_{m})=O(m^{4})$, and (iii) $Z(\mathcal{F}_{m}^{\diamond})=\Omega(2^{m}/\mathrm{poly}(m))$.
Proof 3.16.
Let $X\coloneqq\{x_{1},\ldots,x_{2m}\}$ and $Y\coloneqq\{y_{1},\ldots,y_{m^{2}}\}$. For $k=1,\ldots,m$, we set $S_{k}\coloneqq\{y_{m(k\!-\!1)\!+\!1},y_{m(k\!-\!1)\!+\!2},\ldots,y_{m(k\!-\!1)\!+\!m}\}$, and $S_{m+k}\coloneqq\{y_{k},y_{m\!+\!k},\ldots,y_{m(m\!-\!1)\!+\!k}\}$.
@@
For minimal hitting set operation, we consider a family of subsets of $Y$. We define $\mathcal{F}_{m}\coloneqq\{S_{1},\ldots,S_{2m}\}$. Since the ZDD size can be bounded by the sum of cardinality of a set in the family [16], $Z(\mathcal{F}_{m})\leq\sum_{i}|S_{i}|=O(m^{2})$. For $S\subseteq\{y_{1},\ldots,y_{m^{2}}\}$, we associate a binary $m\times m$ matrix, where the $(i,j)$-element is $1$ if and only if $y_{m(i-1)+j}\in S$. Then, $S\cap S_{k}\neq\emptyset$ means that the $k$-th row of the matrix has at least one $1$ and $S\cap S_{m+k}\neq\emptyset$ means that the $k$-th column of the matrix has at least one $1$. Thus, $S\in\mathcal{F}_{m}^{\sharp}$ if and only if the corresponding matrix has at least one $1$ for any column or row and no proper subset of $S$ satisfies this property. The minimal matrix having this property is the permutation matrix, and thus $\mathcal{F}_{m}^{\sharp}=\mathcal{P}_{m}$, that is, the permutation function. This implies $Z(\mathcal{F}_{m}^{\sharp})=\Omega(2^{m}/m)$.
For closure operation, we consider a family of subsets of $X\cup Y$. For $k=1,\ldots,m$ and $\ell=1,\ldots,m$, we define $R_{k,\ell}\coloneqq(X\setminus\{x_{k},x_{m\!+\!\ell}\})\cup((Y\setminus S_{k}\setminus S_{m\!+\!\ell})\cup\{y_{m(k\!-\!1)\!+\!\ell}\})$. We define $\mathcal{F}_{m}\coloneqq\{R_{k,\ell}\mid k,\ell=1,\ldots,m\}$. Again, since the ZDD size can be bounded by the sum of cardinality of a set in the family [16], $Z(\mathcal{F}_{m})\leq\sum_{k,\ell}|R_{k,\ell}|=O(m^{4})$. Then, we show that $\mathcal{F}_{m}^{\cap}\cap\mathcal{C}_{m}=\mathcal{P}_{m}$, where $\mathcal{P}_{m}$ is the permutation function. If it is shown, $Z(\mathcal{F}_{m}^{\cap}\cap\mathcal{C}_{m})=Z(\mathcal{P}_{m})=\Omega(2^{m}/m)$. On the other hand, $Z(\mathcal{F}_{m}^{\cap}\cap\mathcal{C}_{m})=O(Z(\mathcal{F}_{m}^{\cap})Z(\mathcal{C}_{m}))$. Since $Z(\mathcal{C}_{m})=O(m^{3})$, we can deduce that $Z(\mathcal{F}_{m}^{\cap})=\Omega(2^{m}/\mathrm{poly}(m))$.
We now prove $\mathcal{F}_{m}^{\cap}\cap\mathcal{C}_{m}=\mathcal{P}_{m}$. First, we show that $\mathcal{F}_{m}^{\cap}\cap\mathcal{C}_{m}\subseteq\mathcal{P}_{m}$. $R_{k,\ell}$ does not contain any element in $S_{k}$ and $S_{m\!+\!\ell}$ except for $y_{m(k\!-\!1)\!+\!\ell}$. By fixing $k$, if $\mathcal{F}^{\prime}\subseteq\mathcal{F}$ contains at least one $R_{k,\ell}$ for some $\ell$, $S=\bigcap_{S^{\prime}\in\mathcal{F}^{\prime}}S^{\prime}$ contains at most one element from $S_{k}$. Moreover, $S$ does not contain $x_{k}$ if and only if $\mathcal{F}^{\prime}$ contains at least one $R_{k,\ell}$ for some $\ell$. Similarly, by fixing $\ell$, if $\mathcal{F}^{\prime}$ contains at least one $R_{k,\ell}$ for some $k$, which is equivalent to that $S$ does not contain $x_{m\!+\!\ell}$, $S$ contains at most one element from $S_{m\!+\!\ell}$. Now we can say that when $S$ contains no element in $X$, $S$ contains at most one element in $S_{k}$ for any $k=1,\ldots,2m$. This means that if $S$ contains no element in $X$ and $m$ elements in $Y$, $S\in\mathcal{P}_{m}$. Thus, $\mathcal{F}_{m}^{\cap}\cap\mathcal{C}_{m}\subseteq\mathcal{P}_{m}$. Next, we show that $\mathcal{F}_{m}^{\cap}\cap\mathcal{C}_{m}\supseteq\mathcal{P}_{m}$. Let $\sigma$ be an arbitrary permutation of $1,\ldots,m$. Then, $\{y_{\sigma(1)},y_{m\!+\!\sigma(2)},\ldots,y_{(m-1)m\!+\!\sigma(m)}\}=R_{1,\sigma(1)}\cap R_{2,\sigma(2)}\cap\cdots\cap R_{m,\sigma(m)}$. This means that any set in $\mathcal{P}_{m}$ is in $\mathcal{F}_{m}^{\cap}$. Thus, $\mathcal{F}_{m}^{\cap}\cap\mathcal{C}_{m}\supseteq\mathcal{P}_{m}$. This concludes $\mathcal{F}_{m}^{\cap}\cap\mathcal{C}_{m}=\mathcal{P}_{m}$.